Circular No. 7848

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bq IN MCG –06-21-1 L. A. G. Monard, Pretoria, South Africa, reports his discovery of an

apparent supernova (mag ∼ 16.1) on three unfiltered CCD frames taken with a 0.30-m reflector on Mar. 9.85 UT; an image taken by Monard through clouds on Mar. 10.80 showed the new object at mag ∼ 16.0. SN 2002bq is located at α = 9^{h}30^{m}50^{s}.92, δ = −35^{o}42′06′′.0 (equinox 2000.0), which is in a dense star region, 35′′ west and 46′′ south of the center of MCG –06-21-1 = ESO 373-5. A. Wassilieff, Palmerston North, New Zealand, obtained a CCD image of the new object on Mar. 11 at mag 16.8 and position end figures 50^{s}.88, 05′′.8. Monard notes that nothing is visible at this location on Palomar Sky Survey digitized images or on an image taken by him on Jan. 14.00 (limiting mag 17.0).

SUPERNOVA 2002bo IN NGC 3190 H. Kawakita and K. Kinugasa, Gunma Astronomical Observatory

(GAO); K. Ayani, Bisei Astronomical Observatory; and H. Yamaoka, Kyushu University, report that a low-resolution spectrum (range 380–750 nm) of SN 2002bo (IAUC 7847), obtained on Mar. 9.6 UT with the GAO 0.65-m telescope (+ GCS), shows it to be a type-Ia supernova, ∼ 2 weeks before maximum light. The continuum suggests that it is highly reddened. The expansion velocity derived from the Si II absorption minimum (rest 635 nm) is ∼ 19 000 km/s.

S. Benetti, G. Altavilla, A. Pastorello, M. Riello, and M. Turatto, Osservatorio Astronomico di Padova; E. Cappellaro, Osservatorio Astronomico di Capodimonte; and T. Tomov and M. Mikolajewski, Torun, write: “Preliminary reduction of a spectrum (range 400–780 nm, resolution 2.5 nm) of SN 2002bo, obtained on Mar. 10.04 UT with the Asiago 1.82-m telescope (+ AFOSC), shows that it is a type-Ia supernova, ∼ 10 days before maximum. The spectrum shows the strong P-Cyg profiles of Si II 635- and 596-nm over a relatively red continuum. The spectrum resembles that of SN 1984A obtained 8 days before maximum light. The expansion velocity measured from the Si II absorption is 17 700 km/s.”

SUPERNOVA 2002bm IN MCG –01-32-19 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Center for Astro-

physics, report that a spectrum (range 370–750 nm) of SN 2002bm (cf. IAUC 7845), obtained by V. Hradecky on Mar. 9.30 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-Ic supernova.

2002 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green